Mother is three times as old as her son. Ten years from now, mother will be two times as old as her son.how old are they now?
Please solve this problem involving systems of linear equations in two variables.

Answer by fractalier(6550)

Call the son's age now, x.
Call the mom's age now, y.
y = 3x
In ten years
y+10 = 2(x+10)
Substitute for y in the second equation and get
3x+10 = 2(x+10)
3x + 10 = 2x + 20
x = 10
y = 30